OpenStudy (ujjwal):

Show that \(2450^n-1370^n-1150^n-250^n\) is divisible by 1980 for all \(n\epsilon N\)

OpenStudy (anonymous):

Are you trying to denote the set membership of n in N? If so use \in, also what is N?

OpenStudy (ujjwal):

n belongs to the set of Natural numbers

OpenStudy (anonymous):

You should use \mathbb{N}

OpenStudy (anonymous):

$$n\in \mathbb{N}$$

OpenStudy (anonymous):

Do you know some modular arithmetic?

OpenStudy (ujjwal):

yeah.. I just don't know how to choose the correct figures ..

OpenStudy (anonymous):

Since \(1980=4\cdot 9\cdot 5\cdot 11\), one way to do this is to try and show that the statement is divisible by 4, 9, 5 and 11.

OpenStudy (anonymous):

oops its 1980 my bad

OpenStudy (anonymous):

Reduce 2450 modulo 1980, then if you can show the expression is zero modulo 4,9,5 and 11 your done.

OpenStudy (ujjwal):

can't i directly show that the statement is divisible by 5 and 396 and hence prove that it is also divisible by 1980?

OpenStudy (anonymous):

That works too.

OpenStudy (anonymous):

yes, sense there both coprime

OpenStudy (ujjwal):

Thanks! So, if i come across such problems, all i need to do is find the factors first and then prove that the expression is divisible by a prime factor among them and the product of the rest all factors.. (as long as the product of rest all factors is not a multiple of the prime number itself)

OpenStudy (anonymous):

Yes, but for a proof of this and other techniques I would read up on some elementary number theory, if you plan on using it in the future or somthing.

OpenStudy (ujjwal):

elementary number theory like?

OpenStudy (anonymous):

Modular arithmetic

OpenStudy (ujjwal):

yeah, i know modular arithmetic.. but i didn't know how to apply that in solving this.. Now i do!

OpenStudy (anonymous):

If gcd(a,b)=1, and a divides n, and b divides n. Then ab divides n, this is essentially what were using here, to simplify the problem into cases where the expression is divisible by smaller integers.

OpenStudy (ujjwal):

Ah, yes!! Thanks!! OS is always helpful! :)

OpenStudy (anonymous):

I might be wrong, but I dont think your statement is divisible by 9. Maybe one of those minus signs should be a plus? Then it will work nicely.

OpenStudy (anonymous):

If the problem was \[2450^n-1370^n+1150^n-250^n\]it would make more sense.

OpenStudy (ujjwal):

Ah yes, That was a typo.. Sorry! Its \(+1150^n\)